HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

INFLUENCE OF TRANSLATIONAL AND ROTATIONAL ENERGY ON THE RATE OF ION-MOLECULE REACTIONS NEAR 0 K: D_{2}^{+}+NH_{3} AND

D_{2}^{+}+ND_{3}

\underline{R. HAHN}, V. ZHELYAZKOVA, D. SCHLANDER, F. MERKT, Laboratory of

Physical Chemistry, ETH Zürich, Vladimir-Prelog-Weg 1-5/10, 8093 Zürich, Switzerland

Ion-molecule reactions are important reactions in atmospheric chemistry, astrophysics and plasma physics^{1}. Reliable values for the reaction rates are key to modelling the relevant reaction networks, but are difficult to study in the collision energy range relevant for astrophysics because ions are very easily heated up by stray electric or magnetic fields. To study the reactions D_{2}^{+}+NH_{3} as well as D_{2}^{+}+ND_{3} at low collision energy, we replace D_{2}^{+} by D_{2} in a Rydberg state (of high principal quantum number n ≥ 30). The distant Rydberg electron shields the ion core from external fields without influencing the ion-molecule reaction within its orbit. We use a Rydberg-Stark deflector to merge a beam of D_{2} Rydberg molecules with a supersonic beam of NH_{3} or ND_{3} and tune the collision energy E_{coll} from ≈ k_{B}· 150 mK to ≈ k_{B}· 70 K^{2}. Short (≈ 20 μs) gas pulses ensure a high collision-energy resolution. We observe two reaction channels for the reaction with NH_{3}: NH_{3}^{+}+D_{2} and NH_{2}D^{+}+D, but these two channels are not resolvable by mass spectrometry for D_{2}^{+}+ND_{3}. The rate coefficients deviate strongly from Langevin rates and increase sharply below 2 K, which is explained by rotationally adiabatic capture calculations as arising from the linear Stark effect of the two inversion tunneling components of ammonia in the electric field of the ion ^{3}^{,}^{4}^{,}^{5}. We study the influence of the rotational temperature (14 K or 39 K) of NH_{3} and ND_{3} using pure ammonia beams and beams of ammonia seeded in helium and measure the rotational states populations with (2+1) REMPI spectroscopy. We compare these results with those obtained for the He^{+}+NH_{3}^{6} and Rg^{+}+NH_{3} (Rg=Ar, Kr, Xe) reactions^{7}.

1doi:10.1038/s42254-020-0198-0, B. A. McGuire, O. Asvany, S. Brünken, S. Schlemmer, Nat. Rev.

Phys., 2, 402-410, (2020).

2doi:10.1103/PhysRevA.90.043403, P. Allmendinger, J. Deiglmayr, J. A. Agner, H. Schmutz, F.

Merkt, Phys. Rev. A, 90, 043403, (2014).

3doi:10.1063/1.453701, J. Troe, Chem. Phys., 87, 2773-2780, (1987).

4doi:10.1080/00268978500100461, D. C. Clary, Mol. Phys., 54, 605-618, (1985).

5doi:10.1063/1.472479, J. Troe, J. Chem. Phys., 105, 6249-6262, (1996).

6doi:10.1039/D1CP03116C, V. Zhelyazkova, F. B. V. Martins, J. A. Agner, H. Schmutz, F. Merkt, Phys. Chem. Chem. Phys., 23, 21606, (2021).

7doi:10.1039/D1SC01652K, A. Tsikritea, K. Park, P. Bertier, J. Loreau, T. Softley, B. R. Heazlewood, Chem. Sci., 12, 10005-10013, (2021).

p-number: p006  Submitted on Tue, 23 May 2023 16:26:47 +0200